OpenStudy (anonymous):

are sinx and sin5x both bound by -1 and 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

why? i don't understand

zepdrix (zepdrix):

The period of the function ( the rate at which the function cycles through all of it's values ranging from 1 to -1, twice ), is affected by the 5. To change the magnitude or size of the sine wave, you need a constant coefficient multiplying in front.\[\Large \color{orangered}{A} \sin x, \qquad\qquad\qquad \color{orangered}{B} \sin 5x\] The first function would range from A to -A because it's being multiplied by A while the second function ranges from B to -B.

zepdrix (zepdrix):

So maybe this would make more sense. \[\Large \sin x, \qquad\qquad\qquad \color{orangered}{5}\sin x\]The second function will go up to 5 and down as far as -5 because of the change in magnitude.

zepdrix (zepdrix):